Reports 1-1 of 1 Clear search Modify search
AdV-ISC (Commissioning up to first full interferometer lock)
gosselin, spinicelli, ruggi, capocasa, derossi, chiummo, bersanetti - 20:43 Thursday 05 May 2022 (55725) Print this report
PR centering

After lunch we started to work again on the PR centering. According to the estiamtion there was only 1 mm missing, so we started with a different strategy than the other times. We were moving the PR and were vertivcally shifting SIB1 to recover the alignment in the arms. We did it in two steps to be able to recover the alignment of the green on EIB.
We checked the angular to longitudinal coupling for both the vertical and horizontal axis and it appears to be slightly worse than what it was after the last shift on PR centering (see plot 1). The beam was also miscentered on the inputs.
We decided to move M3 and M4 to recover the centering on the inputs and see the effects on the PR.

gps 1335797245 M3_TY 200steps backward
gps 1335797367 M4_TX 120steps forward
gps 1335797951 M3_TY 330steps backward
gps 1335797986 M4_TX 200steps forward

On the horizontal axis

gps 1335799184 M4_TY 360steps backward
gps 1335799231 M3_TX 250steps backward
gps 1335799880 M4_TY 70steps backward
gps 1335799899 M3_TX 50steps backward

Those steps allow us to reduce the coupling on the inputs. However when it came to measure the effect on PR, the recovery of the lock of the CITF was very difficult (more details in another entry).

Images attached to this report
Comments to this report:
capocasa, ruggi - 23:32 Thursday 05 May 2022 (55726) Print this report

During the beam shift activity, we took the chance to check the newly implemented channels to monitor beam shift on NI/WI.

Dithering lines generated in the Acl_pre process are sent to input marionettes with the following frequencies: WI/NI_TX: 1.5 Hz, WI/NI_TY: 1.3 Hz.

Then B7/B8 are demodulated at the frequency of the lines providing the monitoring signals ASC_{NI,WI}_{TX,TY}_LF_{B7,B8}_{I,Q}_FS

Fig.1 shows such demodulated signals (I and Q), accounting for NI/WI decentering (both horizonal and vertical), during the beam shift.

The demodulation phase has been tuned offline by taking atan(I/Q).

Fig 2. shows the magnitude of the signals. We plan to do a proper calibration of the signals and turn them on on regular basis to monitor the beam shift.

A function to switch on/off the lines has been implemented on

{WI,NI}_LF_line_toggle (ampl_TX,ampl_TY). The amplitude used today was 1e2.

Images attached to this comment
Search Help